Please help me! This entire lesson confuses me!
if m(x)=x+5/+-1 and n(x) = x – 3, which function has the same domain as m(n(x))?

Answer by Cromlix(4381)

Hi there,
Not sure about your +- on bottom
line but :-
if m(x)=x+5/+-1 and n(x) = x – 3,
which function has the same domain as m(n(x))?
m(x) = x + 5/+-1 n(x) = x - 3
m(n(x)) = [(x - 3) + 5]/+- 1
m(n(x)) = [x + 2]/+-1
Hope this helps :-)